# Inequality in a model of capitalist economy

### v1

@@ IV.3 Influence of the wage interval
Figure 6: Distributions of (a) wealth $F(w)$ and (b) income $F(y)$, for different wage ranges $[p_{a},p_{b}]$ indicated in the legend. The bimodal case corresponds to $p=10$ or $p=90$, with equal probability 1/2. $\overline{w}=100$ (blue), 600 (red), and $N=10^{3}$. Note that, in all cases, $\overline{p}=50$, and the results coincide for same $\overline{w}$.
We analyze various wage ranges, $[p_{a},p_{b}]$, where salaries are uniformly distributed, and explore their impact on the distribution of wealth. By varying the wage intervals while keeping the mean value $\overline{p}$ fixed, we found that the results remain notably invariant to the wage range, as shown in Fig. 6(a). The same holds for bimodal distributions, as illustrated in the figure for $p=10$ and 90, with probability 1/2. We also explored non-uniform salary distributions, finding again no significant differences. Consequently, a significant conclusion not addressed in the original model is that the effect of wages is determined by their mean value $\overline{p}$. Furthermore, the same results are consistently observed for the same ratio $R=\overline{w}/\overline{p}$. Thus, decreasing $\overline{p}$ exerts an identical effect as increasing $\overline{w}$, both enhancing inequality.
@@ V Income distribution
In this section, we focus on the distribution of incomes (earnings), applying a similar analysis as for the wealth distribution. The results reveal similar trends in relation to inequality.
Figure 7: Annual income distribution due to workers’ wages and due to the revenue of firms. We used $N=102400$, $\overline{w}=100$, and salary range $[10,90]$. The segment with slope -1 is drawn for comparison.
We consider income to be the sum of all remunerations received by the agent over a one-year period. A worker receives a wage (rule 6), while a capitalist receives a revenue, directly or through its employees (rule 4). Additionally, if an agent changes status within a year, they can receive both types of earnings: wages and revenues.
The separate distributions of worker’s wages and of firms’ revenues are shown in Fig. 7, exhibiting a long tail for revenues. Therefore, the distribution of total income, in Figs. 8 and 9, has two regimes that can be associated to each source of income.
Figure 8: (a) Annual income distribution $F(y)$ of the full population and (b) corresponding Lorenz curves, for different system sizes $N$ indicated in the legend, with $\overline{w}=100$ and wage range $[10,90]$. In (a), the line with slope -1 is drawn for comparison. In (b), the inset shows the average annual Gini coefficient vs $N$.
@@ V.1 Influence of the system size $N$
In Fig. 8 we can see how the distribution of wealth is affected by the size of the system. In panel (a), we can see how the two-regime distribution evolves as we increase the number of agents, with a Pareto regime similar to that of the wealth distribution. In the inset of panel (b), we observe a slow increase of the inequality with $N$, as measured by the Gini coefficient, similarly to the behavior of wealth distribution with $N$.
@@ V.2 Influence of the wealth per capita $\overline{w}$
Fig. 9 shows how the system responds when we increase the wealth per capita. Similarly to the behavior found in the previous section for the wealth distribution, a strong increase of income inequality also occurs, as can be measured by the Gini coefficient, shown in the inset of panel (b), which originates from the cumulative wealth distribution and Lorenz curves displayed in Figs. 9 (a) and (b), respectively.
Figure 9: (a) Complementary cumulative distribution of annual income $F(y)$, and (b) corresponding Lorenz curves, for different values of the wealth per capita. Gini index vs. $\overline{w}$ is in the inset. The system size is $N=1000$ and the wage range is [10,90].
@@ V.3 Influence of wage interval
Regarding the wage interval, we confirm here the same behavior observed in wealth distributions, which is dependent only on the average wage $\overline{p}$, as shown in Fig. 6(b). Therefore, increasing $\overline{p}$ produces a rapid decay of income inequality.
VI Conclusions and final remarks
We have conducted a detailed study of the so-called Social Architecture (SA) model, introduced by Wright in 2005 [2]. The model is based on a few simple rules that govern status changes and wealth exchanges. The former occur through hiring and firing, while wealth exchange happens in two ways: directly, as employers pay salaries, and indirectly, as (all) agents contribute to a common fund from which employers retrieve their profits.
Wright presented the model with four parameters: system size $N$, wealth per capita $\overline{w}$, and the lower/upper limits of the salary range, $p_{a}$ and $p_{b}$. Regarding system size, we have shown that all relevant results of the model are independent of $N$, even for systems as small as $N=100$. As for the salary range, the results are sensitive only to the average value $\overline{p}=(p_{a}+p_{b})/2$.
Moreover, since the quantities in the rules can be scaled—indicating a choice of monetary unit without altering the dynamics—we conclude that the only remaining relevant parameter is the ratio $R=\overline{w}/\overline{p}$. On the one hand, $R\geq 1$ is required for the dynamics to evolve from the initial state, as agents would otherwise not have enough wealth to hire employees on average. On the other hand, if $R$ is too large, it leads to an anomalous concentration of wealth and a freezing of the dynamics. Remarkably, within the regime $R\geq 1$, this single parameter model yields a two-regime distribution of wealth and income: Boltzmann-Gibbs-like for low values and Pareto-like at the tails.
@@ VI Conclusions and final remarks
According to the model’s predictions, increasing total wealth in the system, while keeping wages fixed, does not alleviate inequality. On the contrary, it amplifies wealth concentration within the capitalist class and reduces the average wealth of employees, thereby increasing inequality, as evidenced by the wealth distribution and the Gini index. Similar effects are observed with respect to income. These outcomes align with the argument that unregulated capitalism drives increasing inequality [15]. As we have seen, the only parameter that matters is $R$, therefore, increasing the average wage has the same effect as decreasing the total wealth.
@@
The model yields a high level of mobility between classes, as shown in Fig. 10, where we plot the average uninterrupted time each agent spends in a given class as a function of wealth per capita. For $\overline{w}\simeq 100$, the time spent in the worker and capitalist classes is shorter than one year and increases with $\overline{w}$ becoming larger for workers, while the agents remain less time unemployed on average as $\overline{w}$ increases. This mobility may seem unrealistic at first sight. However, note, for instance, that according to data from the Bureau of Labor Statistics of US, 50% of the private establishments fail in the first five years [16]. According to the model, this can be understood by observing the higher concentration of wealth that makes capitalist agents more robust against decaying to the unemployed class. Let us also note the homogeneity of the agents, in the sense that they are statistically equivalent and will visit equally the different classes along time, except for very large wealth per capita, in which case freezing occurs.
@@
Some assumptions of the model may be revised, leading to future research perspectives. One such assumption is the expenditure rule. In its present form, an agent could potentially spend all of their wealth in a single transaction. This is clearly extreme when considering a real economy. Modifying this rule could help extend the persistence of capitalists within their class. In fact, this issue is crucial for employers, as the model makes no distinction between personal and firm wealth, meaning that a business owner could exhaust both in a single transaction! In addition, employers hire new employees whenever the opportunity arises, without limitations. It could be interesting to impose constraints related to market value, such as limiting spending to employee savings or firm cash reserves.
Another point is that, in this model, agents are homogeneous and indistinguishable apart from the class to which they belong, with no variation in skill levels. Specifically, hiring assumes that unemployed agents are actively seeking jobs, while employed agents are not pursuing market reallocation, even though people may seek salary upgrades. Some variants addressing this aspect have already been explored [4].
Overall, considering that this is a one-parameter model (governed by $R$), the results are remarkable, offering valuable perspectives on the dynamics of wealth distribution and inequality within a simplified economic framework.
Acknowledgments
@@
We acknowledge financial support from from CAPES, grant #88887.969088/2024-00 and finance code 001. C.A. acknowledges partial financial support from Brazilian agencies CNPq (311435/2020-3) and FAPERJ (CNE E-26/201.109/2021). S.G. would also like to thank CNPq for partially supporting this work under grant #314738/2021–5.

### v2

@@ IV.3 Influence of the wage interval
We analyze various wage ranges, $[p_{a},p_{b}]$, where salaries are uniformly distributed, and explore their impact on the distribution of wealth. By varying the wage intervals while keeping the mean value $\overline{p}$ fixed, we found that the results remain notably invariant to the wage range, as shown in Fig. 6(a). This is in agreement with the results for uniform distributions reported in Ref. Lin (2012). The same holds for bimodal distributions, as illustrated in the figure for $p=10$ and 90, with probability 1/2. We also explored non-uniform salary distributions, finding again no significant differences. Consequently, a significant conclusion not addressed in the original model is that the effect of wages is determined by their mean value $\overline{p}$. Furthermore, the same results are consistently observed for the same ratio $R=\overline{w}/\overline{p}$. Thus, decreasing $\overline{p}$ exerts an identical effect as increasing $\overline{w}$, both enhancing inequality.
Figure 6: Complementary cumulative distributions of (a) wealth, $F(w)$, and (b) income, $F(y)$, for different wage ranges $[p_{a},p_{b}]$ indicated in the legend. The bimodal case corresponds to $p=10$ or $p=90$, with equal probability 1/2. $\overline{w}=100$ (blue), 600 (red), and $N=10^{3}$. Note that, in all cases, $\overline{p}=50$.
@@ V Income distribution
In this section, we focus on the distribution of incomes (earnings), applying a similar analysis as for the wealth distribution. The results for income reveal trends similar to those for wealth in relation to inequality. Let us remark that firm owners (capitalists) collect revenues, while workers receive salaries (wages), and both are computed as incomes.
We consider income to be the sum of all remunerations received by the agent over a one-year period. A worker receives a wage (rule 6), while a capitalist receives a revenue, directly or through its employees (rule 4). Additionally, if an agent changes status within a year, they can receive both types of income: wages and revenues.
The separate distributions of worker’s wages and of firms’ revenues are shown in Fig. 7, exhibiting a long tail for revenues. A qualitatively similar result has been previously reported by L. Lin Lin (2012), although for other parameter values. Therefore, the distributions of total income, Fig. 8 (as a function of system size) and Fig. 9 (as a function of wealth per capita), both present two regimes that can be associated to each source of income. We have shown a similar decomposition before, when discussing wealth, as illustrated in Fig. 2.
Figure 7: Annual income cumulative distribution due to workers’ wages and the revenue of firms. We used $N=102400$, $\overline{w}=100$, and salary range $[10,90]$. The segment with slope -1 is drawn for comparison.
@@ V.1 Influence of the system size $N$
In Fig. 8 we show the impact of the size of the system on the distribution of wealth. In panel (a), we can see how the two-regime distribution evolves as we increase the number of agents, with a Pareto regime similar to that of the wealth distribution. In the inset of panel (b), we observe a slow increase of the inequality with $N$, as measured by the Gini and Kolkata indexes, similarly to the behavior of wealth distribution with $N$.
Figure 8: (a) Annual income complementary cumulative distribution $F(y)$ of the full population and (b) corresponding Lorenz curves, for different system sizes $N$ indicated in the legend, with $\overline{w}=100$ and wage range $[10,90]$. In (a), the line with slope -1 is drawn for comparison. In (b), the inset shows the average annual Gini and Kolkata indexes vs $N$.
@@ V.2 Influence of the wealth per capita $\overline{w}$
Fig. 9 shows how the system responds when we increase the wealth per capita. Similarly to the behavior found in the previous section for the wealth distribution, a strong increase of income inequality also occurs, as measured by the Gini and Kolkata indexes, shown in the inset of panel (b), which originates from the cumulative wealth distribution and Lorenz curves displayed in Figs. 9 (a) and (b), respectively.
Figure 9: (a) Annual income complementary cumulative distribution $F(y)$ of the full population $F(y)$, and (b) corresponding Lorenz curves, for different values of the wealth per capita. Annual Gini and Kolkata indexes vs. $\overline{w}$ is in the inset. The system size is $N=1000$ and the wage range is [10,90].
@@ V.3 Influence of wage interval
Regarding the wage interval, we confirm here the same behavior observed in wealth distributions, which depends only on the average wage $\overline{p}$, as shown in Fig. 6(b). Therefore, increasing $\overline{p}$ produces a rapid decay of income inequality.
VI Final remarks and perspectives
We have conducted a detailed study of the so-called Social Architecture (SA) model, introduced by Wright in 2005 Wright (2005). The model is based on a few simple rules that govern status changes and wealth exchanges. The former occur through hiring and firing, while wealth exchange happens in two ways: directly, as employers pay salaries, and indirectly, as (all) agents contribute to a common fund from which employers retrieve their profits. In contrast to most models in the literature discussed in Sec. I, starting from initial conditions in which agents are in equal economic and social positions, the the SA model system self-organizes into two classes with a wealth distribution characterized by the two-regime real-world shape.
Wright presented the model with four parameters: system size $N$, wealth per capita $\overline{w}$, and the lower/upper limits of the salary range, $p_{a}$ and $p_{b}$. Regarding system size, we have shown that for sufficiently large $N$, the average values of wealth and income within each class become independent of $N$, although long-tails develop with increasing $N$ for the capitalist class. Moreover, the same qualitative results are observed even for systems as small as $N=100$. In terms of the salary range, the results are sensitive only to the average value $\overline{p}=(p_{a}+p_{b})/2$. Moreover, since the quantities appearing in the governing rules can be scaled —indicating a choice of monetary unit without altering the dynamics— we arrive to the conclusion that the only remaining relevant parameter is the ratio $R=\overline{w}/\overline{p}$. On the one hand, $R\geq 1$ is required for the dynamics to evolve from the initial state, as agents would otherwise not have enough wealth to hire employees on average. Furthermore, if $R$ is too large, it leads to an anomalous concentration of wealth and a freezing of the dynamics. Remarkably, within the regime $R\geq 1$, this single parameter model yields a two-regime distribution of wealth and income: Boltzmann-Gibbs-like for low values and Pareto-like at the tails.
To quantify the inequality in wealth and income distributions, we utilized both the Gini and the Kolkata indexes Banerjee et al. (2020). Their dependence on the system size $N$ and on the average wealth $\bar{w}$, shown in the insets of Figs. 4b, 5b, 8b and 9b, display similar trends, as expected. According to the model’s predictions, we have seen that increasing total wealth in the system, while keeping wages fixed, does not alleviate inequality. On the contrary, it amplifies wealth concentration within the capitalist class and reduces the average wealth of employees, thereby increasing inequality, as evidenced by the wealth distribution and the inequality indexes. Similar effects are observed with respect to income. These results align with the argument that unregulated capitalism drives increasing inequality Piketty (2013). Since $R$ is the relevant parameter, increasing the average wage has the same effect as decreasing total wealth.
@@ VI Final remarks and perspectives
Let us also include and comment an additional simulation result. The model yields a high level of mobility between classes, as illustrated in Fig. 10, where we plot the average uninterrupted time each agent spends in a given class as a function of wealth per capita. For $\overline{w}\simeq 100$, the time spent in the worker and capitalist classes is shorter than one year and increases with $\overline{w}$ becoming larger for workers, while the agents remain less time unemployed on average as $\overline{w}$ increases. This mobility may seem unrealistic at first sight. However, note, for example, that according to data from the US Bureau of Labor Statistics, 50% of private establishments fail in the first five years BLS-US . According to the model, this effect can be understood by observing the higher concentration of wealth that makes capitalist agents more robust against decaying to the unemployed class.
@@
Let us now discuss some model assumptions that could be revised, potentially opening new research perspectives. One such assumption is the expenditure rule. In its current form, an agent could theoretically spend their entire wealth in a single transaction. This is clearly extreme when considering a real economy. Modifying this rule could help extend the persistence of capitalists within their class. In fact, this issue is crucial for employers, as the model makes no distinction between personal and firm wealth, meaning that a business owner could exhaust both in a single transaction! In addition, employers hire new employees whenever the opportunity arises, without limitations. It could be interesting to impose constraints related to market value, such as limiting spending to employee savings or firm cash reserves. Another point is that, in this model, agents are homogeneous and indistinguishable apart from the class to which they belong, with no variation in skill levels. Specifically, hiring assumes that unemployed agents are actively seeking jobs, while employed agents are not pursuing market reallocation, even though people may seek salary upgrades. Some variants addressing this aspect have already been explored Wright (2009).
Overall, considering that, as shown in the present work, this is essentially a one-parameter model (namely, $R$), the results are remarkable. Unlike other valuable models which reproduce the Pareto law or even the two regimes mentioned in Section I, the SA model stands out for its simplicity. It requires no additional rules or interventions to produce a two-class distribution, offering a fundamental approach for further investigation on the dynamics of wealth distribution and inequality within a simplified economic framework.
@@
Acknowledgments: We acknowledge financial support from CAPES, grant 88887.969088/2024-00 and finance code 001. C.A. acknowledges partial financial support from Brazilian agencies CNPq (311435/2020-3) and FAPERJ (CNE E-26/204.130/2024). S.G. would also like to thank CNPq for partially supporting this work under grant 314738/2021–5.